Consider an oligopoly market with two identical firms. Assume each firm produces with 𝐢(π‘ž)=10π‘ž

C(q)=10q, and market demand curve is 𝑝=50βˆ’2𝑄

p=50βˆ’2Q.


Solve for the Cournot Nash equilibrium quantities, prices, and profits for the two firms. What is the equilibrium quantity each firm produces?γ€€Round off to two decimal places.Β 


1
Expert's answer
2021-04-23T07:39:24-0400

Solution:

Derive Total Revenue (TR):

P = 50 – 2Q

Q = "q_{1} +q_{2}"


Solve for firm 1:

TR1 = P*Q

TR1 = (50 – 2("q_{1} +q_{2} )" ) q1

TR1 = 50q1 – 2q12 - 2q1q2


Derive marginal revenue:

MR1 = derivative of TR with respect to Q

"\frac{\partial TR_{1} }{\partial q_{1} }"Β = 50 – 4q1 - 2q2

MR1 = 50 – 4q1 - 2q2

Compute the profit-maximizing output by setting MR = MC:

MC1 = derivative of TC1 with respect to q1

TC1 = 10q


MC1 ="\frac{\partial TC_{1} }{\partial q_{1} }"Β = 10


MR1 = MC1

50 – 4q1 - 2q2 = 10

50 - 10 - 2q2 = 4q1

40 - 2q2 = 4q1

q1 = 10 - 0.5q2


Solve for firm 2:

TR2 = P*Q

TR2 = (50 – 2("q_{1} +q_{2} )" ) q2

TR2 = 50q2 – 2q22 - 2q1q2


Derive marginal revenue:

MR2 = derivative of TR2 with respect to q2


"\frac{\partial TR_{2} }{\partial q_{2} }"Β = 50 – 4q2 - 2q1


MR2 = 50 – 4q2 - 2q1

Compute the profit-maximizing output by setting MR2 = MC2:

MC2 = derivative of TC2 with respect to q2

TC2 = 10q


MC2 ="\frac{\partial TC_{2} }{\partial q_{2} }"Β = 10


MR2 = MC2

50 – 4q2 - 2q1 = 10

50 - 10 - 2q1 = 4q2

40 - 2q1 = 4q2

q2 = 10 - 0.5q1


Substitute firm 1 into the reaction function of firm 2

q1 = 10 - 0.5q2

q1 = 10 - 0.5(10 - 0.5q1)

q1 = 10 - 5 + 0.25q1

q1 - 0.25q1 = 5

0.75q1 = 5


q1 = "\frac{5}{0.75} = 6.67"


q2 = 10 - 0.5q1

q2 = 10 - 0.5(6.67)

q2 = 6.67

Each firm produces an equilibrium quantity of 6.67


Substitute to derive equilibrium price:

P = 50 - 2Q

P = 50 - 2 (q1 + q2)

P = 50 - 2(6.67 + 6.67)

P = 50 - 26.68

P = 23.32

The market equilibrium price is 23.32


Profit for each firm:

Profit = TR - TC

= "(P\times Q) - TC"

= "(23.32\times 6.67) - 10"

= "155.54 - 10 = 145.54"

Profit = 145.54

Each firm will have a profit of 145.54